OpenStudy (anonymous):

If R the total resistance for a parrallel cicuit with two resistors of resistance r1 and r2, then 1/R=1/r1+1/r2. Find the resistance, r1, if the total resistance is 20 ohms, and the r2 is 75 ohms. Round if necessary.

OpenStudy (anonymous):

You just need to sub in and solve for r1

OpenStudy (anonymous):

1/20 = 1/r1 + 1/75

OpenStudy (anonymous):

so then 1/-55 = 1/x

OpenStudy (anonymous):

No, you don't add fractions like that :P

OpenStudy (anonymous):

I see what you did. Bad! You have to remember how to work with fractions.

OpenStudy (anonymous):

1/20 - 1/75 = 1/r1

OpenStudy (anonymous):

No worries, everyone gets rusty :P. (1*75)/(20*75) - (1*20)/(20*75) = 1/r1

OpenStudy (anonymous):

sorry here are the answers. A. 16 B. 1405 C. 27 D. 102

OpenStudy (anonymous):

they have to have the same denominator in order to add or subtract them. So we just multiply by a factor of one. (1/20 * 75/75)

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

I'm not gonna do it for ya brah. I gave you the answer essentially, give it a try. You won't learn anything if I give you the answer on a silver platter ;).

OpenStudy (anonymous):

I'll let you know if you're wrong

OpenStudy (anonymous):

150 is LCD

OpenStudy (anonymous):

sorry 200

OpenStudy (anonymous):

10/200 - 3/200 = 1r1

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

It's not too important to find the LCD, as long as they both have the same denominator I just did (20*75) to make it easy... in that case I would get 1500 as the Common Denominator.

OpenStudy (anonymous):

Look again at what I wrote here: (1*75)/(20*75) - (1*20)/(20*75) = 1/r1

OpenStudy (anonymous):

Alternately, if you hate fractions like the plague... you can just switch them to decimals. Use your calculator and find out what 1/20 is, then find out what 1/75 is. Then subtract the two.

OpenStudy (anonymous):

1/20 = .05

OpenStudy (anonymous):

1/20 - 1/75 = 1/r1 (1*75)/(20*75) - (1*20)/(20*75) = 1/r1 (75 - 20)/1500 = 1/r1 -55/1500 = 1/r1

OpenStudy (anonymous):

1/75 =.02

OpenStudy (anonymous):

1/20 - 1/75 = 1/r1 (1*75)/(20*75) - (1*20)/(20*75) = 1/r1 (75 - 20)/1500 = 1/r1 55/1500 = 1/r1 r1 = 1500/55 Done

OpenStudy (anonymous):

and that is C thanks so much

OpenStudy (anonymous):

Work on your fractions man. It's going to be very useful in the parallel circuit stuff of electricity.

OpenStudy (anonymous):

Ok will do thanks!!!

OpenStudy (anonymous):

I'm glad to help.